OpenStudy (anonymous):

Use Heisenberg's Uncertainty Principle to calculate the uncertainty in the position of a 546 mg ball moving at a speed of 5.3 m/s if the speed is known to be correct within ±0.05 m/s.

OpenStudy (anonymous):

My Textbook defines the formula as \[\Delta x \times \Delta (mv) \ge \frac{ h }{ 4\pi }\]

OpenStudy (jfraser):

solve for \(\Delta x\)

OpenStudy (anonymous):

Alright, but what should I plug in for each variable?

OpenStudy (jfraser):

every value is given except for \(\Delta x\). Mass is given, and the uncertainty is given, h is a constant, and \(4\pi\) is a constant